calulate the # of generation, generation time and growth rate for the following e coli

bactetia logarthmic
time cell #
9am 100,000
2pm. 100,000,000 cells

ASAP
Answer #1
  • Total time = t= 5 hours = 5 x 60 =300 minutes
  • Initial number of cells = N1=100,000
  • Final number of cells = N2 = 100,000,000
  • Number of generation = n = (log N2 – log N1)/ log 2
  • Number of generation = n = (log 100,000,000- log 100,000)/ 0.301 = (8-5) / 0.301 = 3/0.301 =9.97 = 10 (approx).
  • Generation time= G = t/n = 300/10 =30 minutes = 2 hours
  • Growth rate = K = 1/G = ½ = 2 doubling per hour.